OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x - 2y = -1 2x - y = 4

OpenStudy (anonymous):

x=-1+2y 2x-y=4 sub first equation into second 2(-1+2y)-y=4 -2+4y-y=4 3y=6 y=2 sub y=2, x=-1+2(2) =-1+4 =3 (3,2)